OpenStudy (anonymous):

A magic number is a number that'll divide evenly every number which ends in itself, for instance, 2 is magic because 2 divides into every number finishing in 2(2,12,22,102 etc). Find a rule to determine whether a number is magical.

OpenStudy (anonymous):

I'll give medals for every attempt.

OpenStudy (anonymous):

if you want a guess i would say 2, 5, or any power of 5 or 10?

OpenStudy (anonymous):

lets start with easy ones. 2 works 5 works 10 works

OpenStudy (anonymous):

You're getting hot satellite73.

OpenStudy (anonymous):

25 works

OpenStudy (anonymous):

Yes, now, try to factor what you found. See if patterns emerge.

OpenStudy (anonymous):

so i looks like powers of 5 or powers of 2*5

OpenStudy (anonymous):

Yes, getting better and better.

OpenStudy (anonymous):

2^n * 5*n

OpenStudy (anonymous):

Almost there james. The powers don't have to be necessarily the same.

OpenStudy (anonymous):

2^n * 5^n i meant :)

OpenStudy (anonymous):

well not exactly. 2^n5^n=10^n so not all

OpenStudy (anonymous):

2^p * 5^q

OpenStudy (anonymous):

maybe 2^n5^m

OpenStudy (anonymous):

There you are. :)

OpenStudy (anonymous):

that looks good to me

OpenStudy (anonymous):

Now some sort of proof, doesn't have to be rigorous...

OpenStudy (anonymous):

satellite did all the work for me...i just tried to capitalize on his work

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

guess my guess was a good guess, even if i guess it was just a guess

OpenStudy (anonymous):

must prove that 2^n * 5^n mod n = 0

OpenStudy (anonymous):

Lol.

OpenStudy (anonymous):

exactly like saying that if you divide by a such a number you get a terminating decimal

OpenStudy (anonymous):

james: generalize the numbers, as in, finishing in 5 for instance works because if a number is abc...10 you can make it abc...00 + 10 which factors in 5 ( abc.... + 1).

OpenStudy (anonymous):

finishing in 10 for instance I meant.

OpenStudy (anonymous):

abc...10 = a * 10^p + b * 10^(p-1) + ... + 10 so if you divide that by 10, you get: a * 10^(p-1) + b * 10^(p-2) + ... + 1

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

which is still an integer

OpenStudy (anonymous):

Right. Now guys, expand the generalization to 2^n * 5^m.

OpenStudy (anonymous):

(2^n * 5^m) = (abc...10) * (def...5) on the right track?

OpenStudy (anonymous):

wait...it's (abc...2) * (def...5) ?

OpenStudy (anonymous):

Yep. Kind of. Done. Do you want another challenge? :)

OpenStudy (anonymous):

(abc...2) * (def...5) will be a huge ugly series, but the constant term will be 10, which is divisible by both 5 and 2

OpenStudy (anonymous):

i am not very good at number theory! but that was an interesting problem :)

OpenStudy (anonymous):

:) Glad you liked it. Shame satellite went to other stuff... Well, ok, here's a challenge, something not for now, but to think over and then come back ok? May I? :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

If you tell a person to choose a number smaller than 60 (but positive), and then ask them to 1) divide it by 3 and tell you the remainder, 2) divide the remainder by 4 and tell you the remainder of that, 3) divide that remainder by 5 and tell the remainder of that. Then, the number choosen is 40 * the first remainder + 45 * the second remainder + 36 * the third remainder mod 60.

OpenStudy (amistre64):

a perfect number is any number whose factors ...add up to itself right?

OpenStudy (anonymous):

I don't know? :(

OpenStudy (anonymous):

yes, that's a perfect number. ana described magic numbers.

OpenStudy (anonymous):

perfect numbers = 6, 28 etc.

OpenStudy (anonymous):

Ana, are we supposed to prove that?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

If you want to see more about it, see: "Chinese Remainder Theorem".

OpenStudy (anonymous):

you must be a math professor. you are smart :)

OpenStudy (anonymous):

No. I'm just a student. But I'll take the compliment. hehe. You're smart too.

OpenStudy (anonymous):

Enter the chat.